OpenStudy (yacoub1993):

If the force acting on an object stays the same, then the acceleration of the object is inversely proportional to its mass. If an object with a mass of 9 kilograms accelerates at a rate of per second per second by a force, find the rate of acceleration of an object with a mass of that is pulled by the same force. A. 3 meters per second per second B. 27 meters per second per second C. 18 meters per second per second D. 24 meters per second per second

OpenStudy (yacoub1993):

C......?

OpenStudy (anonymous):

is there any more info for this? i'm a bit unsure on this part, "find the rate of acceleration of an object with a mass of that is pulled by the same force."

OpenStudy (yacoub1993):

no

OpenStudy (yacoub1993):

continue....

OpenStudy (yacoub1993):

@DemolisionWolf continue solving Please

OpenStudy (anonymous):

@Yacoub1993 i'm still trying to figure out how to get a value... I have 1 equation with 3 unknowns. a=(m)/(9a)

OpenStudy (yacoub1993):

ohh ok

OpenStudy (anonymous):

we know the solution will be a multiple of 9, so either 3, 18, 27...

OpenStudy (yacoub1993):

ok which one do u think is correct

OpenStudy (anonymous):

do we know if the 'mass' is larger or smaller than 9kg?

OpenStudy (anonymous):

I think there's a missing info of a given definite acceleration. Lacks info.

OpenStudy (yacoub1993):

Question 14 of 20 5.0 Points If the force acting on an object stays the same, then the acceleration of the object is inversely proportional to its mass. If an object with a mass of 9 kilograms accelerates at a rate of per second per second by a force, find the rate of acceleration of an object with a mass of that is pulled by the same force. A. 3 meters per second per second B. 27 meters per second per second C. 18 meters per second per second D. 24 meters per second per second

OpenStudy (anonymous):

what was question 13 of 20, maybe there is information we should have from the prior question

OpenStudy (yacoub1993):

Question 13 of 20 5.0 Points Find a rational zero of the polynomial function and use it to find all the zeros of the function. f(x) = x4 + 3x3 - 5x2 - 9x - 2 A. {1, -2, -2 + sqrt3 , -2 - sqrt3 } B. {-1, -2, -2 + sqrt5 , -2 - sqrt5 } C. {-1, 3, -2 + sqrt5 , -2 - sqrt5 } D. {-1, 2, -2 + sqrt3 , -2 - sqrt3 }

OpenStudy (yacoub1993):

that is question 13

OpenStudy (anonymous):

" If an object with a mass of 9 kilograms accelerates at a rate of per second per second by a force" F_1=9*a_1 "find the rate of acceleration of an object with a mass of" a_2=F_2 / m_2 "that is pulled by the same force" F_1=F_2 "If the force acting on an object stays the same, then the acceleration of the object is inversely proportional to its mass." a=1/m

OpenStudy (anonymous):

...this sounds like an exam or a quiz then...

OpenStudy (yacoub1993):

quiz

OpenStudy (anonymous):

i'm stumped, the best I can end up with is that solution acceleration = (9*a_1)/m_2 is this a physics class or a math class?

OpenStudy (yacoub1993):

math

OpenStudy (yacoub1993):

precalculas

OpenStudy (anonymous):

hmmmm..

OpenStudy (anonymous):

I'd say it's 'A' only because  and the value can't be greater than 9, because if the value were greater than 9, we would need more information about the second mass. so all we have is 3, 27, 18, 24. ... I'd guess 3